OpenStudy (anonymous):

if f(x)=3/x+2- the square root of x-3 complete the following statement (round your answer to the nearest hundredth): f(7)=

OpenStudy (anonymous):

\[f(x)=3/(x+2)(\sqrt{x-3}); f(7)=3/(7+2)(\sqrt{7-3})=3/(9+\sqrt{4})=3/11=0,272727 \approx0,27\]

OpenStudy (anonymous):

\[0,2727272727\approx0,27\]